the equation y=-1777x+27,153 can be used to predict the number y of gun deaths in the united states, x years after 2000 that is, x= 0 corresponds to 2000, x=3 corresponds to 2003,x=6 corresponds to 2006 and so on.predict the number of gun deaths in 2006 and 2007. in what year will the number of gun deaths be 11,160?
Answer by orca(409)

SOLUTION:
The number of gun deaths y in 2006 can be found by substituting x = 6 into the formula y=-1,777x+27,153.
y = -1,777*6 + 27,153 = 16,491
The number of gun deaths y in 2006 can be found by substituting x = 7 into the formula y=-1,777x+27,153.
y = -1,777*7 + 27,153 = 14,714
To find in what year the number of gun deaths will be 11,160, solve the following equation for x:
-1,777x+27,153 = 11,160
1777x = 15993
x = 9
So in 2009 the number of gun deaths will be 11,160.